# MSR-Codec: A Low-Bitrate Multi-Stream Residual Codec for High-Fidelity Speech Generation with Information Disentanglement

### v2

@@ Abstract
Audio codecs are a critical component of modern speech generation systems. This paper introduces a low-bitrate, multi-scale residual codec that encodes speech into four distinct streams: semantic, timbre, prosody, and residual. This architecture achieves high-fidelity speech reconstruction at competitive low bitrates while demonstrating an inherent ability for information disentanglement. We construct a two-stage language model for text-to-speech (TTS) synthesis using this codec, which, despite its lightweight design and minimal data requirements, achieves a state-of-the-art Word Error Rate (WER) and superior speaker similarity compared to several larger models. Furthermore, the codec’s design proves highly effective for voice conversion, enabling independent manipulation of speaker timbre and prosody. Our inference code, pre-trained models, and audio samples are available at https://github.com/herbertLJY/MSRCodec.
Index Terms—  audio codec, speech generation, low-bitrate, information disentanglement
@@ 1 Introduction
With the fast development of large language models (LLMs), LLM-based systems have become the mainstream approaches in the task of text-to-speech (TTS) and achieved state-of-the-art performance. Central to these systems is the neural audio codec [1, 2, 3, 4], a critical module that transforms continuous speech waveforms into sequences of discrete tokens. In a typical generation pipeline, an LLM predicts these speech tokens conditioned on input text, and a corresponding decoder then synthesizes the final audio from this token sequence.
@@
A primary goal for modern neural codecs is to achieve high compression rates without sacrificing quality. Models like SoundStream [5] and Encodec [6] use stacked convolutions to compress speech by over 100 times, using Residual Vector Quantization (RVQ) to maintain high-fidelity reconstruction. However, this fidelity often comes at the cost of a high bitrate—the amount of data required to store the tokenized speech. For instance, popular codecs may require bitrates of 6kbps or more, which increases the complexity for storage, transmission, and the downstream language models that must predict these longer token sequences.
@@
To address this efficiency challenge, researchers have explored several information disentanglement strategies [7, 8]. One prominent approach separates speech into time-variant frame-level tokens and time-invariant codes that capture global information, such as speaker timbre and environmental sounds. As this global information does not scale with the speech’s length, this method can significantly reduce the required bitrate. For example, LSCodec [9] effectively disentangles timbre by explicitly encoding speaker information with a prompt network. A more structured approach, seen in models like SemantiCodec [10], factorizes speech into its fundamental components: semantic information (linguistic content) and acoustic information. Semantic tokens, which can be extracted via supervised ASR tasks [11] or self-supervised models like Hubert [12], capture what is being said but are insufficient on their own for reconstruction. They must be complemented by another set of acoustic tokens or a generative model to synthesize the final high-fidelity speech.
@@
In this work, we propose a highly efficient, low-bitrate codec that advances this disentanglement paradigm for speech generation. Inspired by factorized representations, we encode speech into four distinct streams: semantic, timbre, prosody, and a fine-grained residual stream. We leverage pre-trained models to extract semantic (Hubert) and timbre (speaker embeddings) information. Our core novelty lies in how these streams are integrated. While prior work [13] relies on complex mechanisms like adversarial training to enforce disentanglement, our model achieves this implicitly through a simple and stable cascaded architecture. Information from each stream is progressively fused via residual connections, naturally encouraging each component to capture distinct attributes of the speech signal. By employing multi-scale processing and smaller codebooks, our codec achieves a compression factor of over 200x (62.5 tokens per second) while maintaining high-quality output. We construct a two-stage LLM for TTS synthesis using our codec, demonstrating that this disentangled representation leads to superior performance. Our lightweight system achieves a state-of-the-art Word Error Rate (WER) and superior speaker similarity. Notably, it outperforms several larger, state-of-the-art models while being trained on a fraction of the data and generating speech significantly faster. Furthermore, the codec’s inherent disentanglement enables highly effective zero-shot voice conversion of both timbre and prosody, confirming the successful separation of these key speech factors.
@@
Fig. 1: The overview structure of the codec, the output frame rate of each module is presented in the figure
@@ 2.1.1 Base Layer: Timbre and Semantic Streams
The foundation of our representation is formed by the speaker’s identity and the linguistic contents.
Timbre Information: We use a pre-trained CAM++ [14] network to extract a single, time-invariant speaker embedding for each utterance, which is $l2$-normalized before feeding into the codec.
@@
The semantic tokens and speaker embedding are fused together using Dec1. The resulting features are downsampled to a 12.5Hz sequence, forming a base representation of the speech’s content and speaker identity. Both the speaker and Hubert encoders remain frozen during training.
@@ 2.1.3 Residual Stream
Acoustic Detail Capture: This stream utilizes the 25Hz feature map produced by Enc1. The residual information is calculated as the difference between Enc1’s output and the upsampled output from the prosody-enhanced stream (from Dec2). This difference is then quantized using a second codebook (VQ2) to create residual tokens.
@@ 2.3 The Structure of the TTS Model
Fig. 2: The network structure of the TTS model. The index denotes the order of each token. Best viewed in colors
@@
Autoregressive Generation: The generation process proceeds autoregressively at 12.5Hz. The input to the next step of the primary Semantic Decoder is formed by concatenating its own output feature from the current step with the embeddings of all tokens generated in that step (semantic, prosody, and residual). Generation terminates when a special stop token is predicted by the Semantic Decoder. The speaker embedding is not used during token prediction; it is only provided as a condition to the codec’s decoder during the final synthesis stage.
@@ 3.1 Datasets and Preprocessing
For training the audio codec, a combination of open-source datasets was used to ensure diversity in speaker identity and prosody. This included the English portion of the Multilingual LibriSpeech (MLS) dataset [15], Textrolspeech [16] for its variety of emotional speech, and VoxCeleb 1 & 2  [17] for its large number of speakers. The TTS model was trained on MLS-en, Libritts [18], and VCTK [19].
All speech audio was resampled to 16kHz. During training, 6-second segments were randomly cropped from each sample. These segments were then transformed into 80-dimensional Mel-spectrograms using a 25ms window length and a 10ms hop length. The codec’s reconstruction ability was evaluated on the Librispeech test set, while the zero-shot TTS capability was evaluated on the English test set from Seed-TTS-eval [^1] https://github.com/BytedanceSpeech/seed-tts-eval.
@@ 3.2.1 Codec Configuration
The encoders (Enc1 &\& Enc2) and decoders (Dec2 &\& Dec3) in the codec are convolution-based networks. Their structures are similar to SEANet [20], with convolution layer parameters adjusted to match the required frame rates for each stream. Two self-attention[21] modules are added at the end of the Enc1 and the input of Dec3, respectively, which involve the global relation between frames. Dec1 is a 2-layer conformer network with cross-attention modules, a structure similar to the decoder in CTX-vec2wav [22]. A pretrained 16kHz Fre-GAN  [23] was used as the vocoder to convert reconstructed Mel-spectrograms into audio waveforms. Two versions of the codec were trained with different bitrates:
MSR-Codec-424: This smaller version uses a codebook of size 32 for the prosody stream (VQ1) and 64 for the residual stream (VQ2), resulting in a bitrate of approximately 424 bps.
MSR-Codec-524: This larger version uses a codebook of size 128 for both the prosody and residual streams, for a bitrate of approximately 524 bps.
For both versions, the semantic stream uses a fixed codebook of 500 centers from the pre-trained Hubert model [^2] https://github.com/facebookresearch/textlesslib/blob/main/textless.
@@ 3.2.3 Evaluation Metrics
  • Reconstruction Quality: Measured with Short-Time Objective Intelligibility (STOI) [24], Perceptual Evaluation of Speech Quality (PESQ) [25], and a pre-trained UTMOS [26] model for naturalness.
  • Speaker Similarity (SIM): Measured between the original and synthesized speech using a WavLM-large-based speaker verification model [27].
@@ 3.3 Results
Table 1: Comparison of speech reconstruction quality on the Librispeech test set. Our proposed models are benchmarked against a comprehensive set of low-bitrate codecs. ‘Nq’ denotes the number of quantizers and ‘TPS’ is tokens per second. Our models’ highest speaker similarity (SIM) and competitive naturalness (UTMOS) scores are in bold.
@@
Table 2: Zero-shot TTS performance on the Seed-TTS English test set. ‘Data size’ is measured in hours. ‘RTF’ denotes Real-Time Factor. Our model’s results are bolded to highlight competitive performance from a much smaller model. The best overall score in each column is also marked.
@@
Table 3: Voice conversion results demonstrating information disentanglement. ‘+S’ indicates timbre conversion and ‘+P’ indicates prosody conversion. Key metrics supporting successful and disentangled transfer are in bold.
@@ 3.3.1 Speech Reconstruction Quality
As shown in  Table 1, the proposed codecs achieve strong reconstruction quality at low bitrates. Compared to other codecs, our models demonstrate the highest speaker similarity and comparable UTMOS scores, indicating that timbre and naturalness are well-preserved. The STOI and PESQ scores are relatively lower, which is expected as these metrics are sensitive to signal-level details that require more bits to reconstruct perfectly. Notably, these scores improve significantly when using the larger bitrate version of the codec.
@@ 3.3.2 Zero-Shot Text-to-Speech
The TTS results in Table 2 highlight our model’s exceptional efficiency and performance. As the most lightweight model trained on the least amount of data, our system still achieves a state-of-the-art Word Error Rate (WER), surpassing all competing models. It also delivers the highest speaker similarity, indicating superior voice cloning. Crucially, our model demonstrates its speed by achieving the lowest Real-Time Factor (RTF), making it the fastest generator among the tested systems. This combination of top-tier accuracy and unparalleled efficiency underscores the significant advantages of our proposed architecture.
@@ 3.3.3 Voice and Prosody Conversion
The codec’s disentangled representation enables inherent voice conversion (VC) capabilities, which we evaluated to validate the separation of speech attributes. $8$ utterances are randomly sampled from the dataset of VCTK. These utterances come from different speakers (4 male/ 4 female), and they are utilized as the target prompts. $100$ utterances are randomly sampled from the clean test set of Libritts and utilized as the source speeches.
Timbre Conversion: By swapping only the speaker embedding of a source speech with a target one, our method achieves higher target speaker similarity than both CosyVoice2 and Seed-VC [28]. As shown in  Table 3, this conversion preserves the original prosody, with the average $F_{0}$ remaining close to the source speech (low $\Delta F_{0,\text{src}}$) and not shifting towards the target (high $\Delta F_{0,\text{tar}}$ ). This confirms the successful disentanglement of the timbre from the prosody.
Prosody Conversion: By keeping the source speaker embedding and semantic tokens but predicting new prosody and residual tokens from a target speech, our model achieves a clear prosody transfer. The results show a significant $F_{0}$ shift towards the target (low $\Delta F_{0,\text{tar}}$) while maintaining the original speaker’s timbre.
Combined Conversion: When both the speaker embedding is swapped and the prosody tokens are re-predicted, the model achieves the best results in transferring both timbre and prosody simultaneously.

### v3

# Li Zhang Ye Guo

@@ Abstract
Audio codecs are a critical component of modern speech generation systems. This paper introduces a low-bitrate, multi-scale residual codec that encodes speech into four distinct streams: semantic, timbre, prosody, and residual. This architecture achieves high-fidelity speech reconstruction at competitive low bitrates while demonstrating an inherent ability for information disentanglement. We construct a two-stage language model for text-to-speech (TTS) synthesis using this codec, which, despite its lightweight design and minimal data requirements, achieves a state-of-the-art Word Error Rate (WER) and superior speaker similarity compared to several larger models. Furthermore, the codec's design proves highly effective for voice conversion, enabling independent manipulation of speaker timbre and prosody. Our inference code, pre-trained models, and audio samples are available at https://github.com/herbertLJY/MSRCodec.
keywords:
audio codec, speech generation, low-bitrate, information disentanglement
@@ 1 Introduction
With the fast development of large language models (LLMs), LLM-based systems have become the mainstream approaches in the task of text-to-speech (TTS) and achieved state-of-the-art performance. Central to these systems is the neural audio codec [wang2023neural, chen2024vall, du2024cosyvoice2, ye2025llasa, guo2024fireredtts], a critical module that transforms continuous speech waveforms into sequences of discrete tokens. In a typical generation pipeline, an LLM predicts these speech tokens conditioned on input text, and a corresponding decoder then synthesizes the final audio from this token sequence.
@@
A primary goal for modern neural codecs is to achieve high compression rates without sacrificing quality. Models like SoundStream [zeghidour2021soundstream] and Encodec [DBLP:journals/tmlr/DefossezCSA23] use stacked convolutions to compress speech by over 100 times, using Residual Vector Quantization (RVQ) to maintain high-fidelity reconstruction. However, this fidelity often comes at the cost of a high bitrate—the amount of data required to store the tokenized speech. For instance, popular codecs may require bitrates of 6kbps or more, which increases the complexity for storage, transmission, and the downstream language models that must predict these longer token sequences.
@@
To address this efficiency challenge, researchers have explored several information disentanglement strategies [ren2024fewer, li2024single]. One prominent approach separates speech into time-variant frame-level tokens and time-invariant codes that capture global information, such as speaker timbre and environmental sounds. As this global information does not scale with the speech's length, this method can significantly reduce the required bitrate. For example, LSCodec [guo25_interspeech] effectively disentangles timbre by explicitly encoding speaker information with a prompt network. A more structured approach, seen in models like SemantiCodec [liu2024semanticodec], factorizes speech into its fundamental components: semantic information (linguistic content) and acoustic information. Semantic tokens, which can be extracted via supervised ASR tasks [du2024cosyvoice] or self-supervised models like Hubert [hsu2021hubert], capture what is being said but are insufficient on their own for reconstruction. They must be complemented by another set of acoustic tokens or a generative model to synthesize the final high-fidelity speech.
@@
In this work, we propose a highly efficient, low-bitrate codec that advances this disentanglement paradigm for speech generation. Inspired by factorized representations, we encode speech into four distinct streams: semantic, timbre, prosody, and a fine-grained residual stream. We leverage pre-trained models to extract semantic (Hubert) and timbre (speaker embeddings) information. Our core novelty lies in how these streams are integrated. While prior work [ju2024naturalspeech] relies on complex mechanisms like adversarial training to enforce disentanglement, our model achieves this implicitly through a simple and stable cascaded architecture. Information from each stream is progressively fused via residual connections, naturally encouraging each component to capture distinct attributes of the speech signal. By employing multi-scale processing and smaller codebooks, our codec achieves a compression factor of over 200x (62.5 tokens per second) while maintaining high-quality output. We construct a two-stage LLM for TTS synthesis using our codec, demonstrating that this disentangled representation leads to superior performance. Our lightweight system achieves a state-of-the-art Word Error Rate (WER) and superior speaker similarity. Notably, it outperforms several larger, state-of-the-art models while being trained on a fraction of the data and generating speech significantly faster. Furthermore, the codec's inherent disentanglement enables highly effective zero-shot voice conversion of both timbre and prosody, confirming the successful separation of these key speech factors.
@@
Figure 1: The overview structure of the codec, the output frame rate of each module is presented in the figure
@@ 2.1.1 Base Layer: Timbre and Semantic Streams
The foundation of our representation is formed by the speaker's identity and the linguistic contents.
Timbre Information: We use a pre-trained CAM++ [wang23ha_interspeech] network to extract a single, time-invariant speaker embedding for each utterance, which is $l2$-normalized before feeding into the codec.
@@
The semantic tokens and speaker embedding are fused together using Dec1. The resulting features are downsampled to a 12.5Hz sequence, forming a base representation of the speech's content and speaker identity. Both the speaker and Hubert encoders remain frozen during training.
@@ 2.1.3 Residual Stream
Acoustic Detail Capture: This stream utilizes the 25Hz feature map produced by Enc1. The residual information is calculated as the difference between Enc1's output and the upsampled output from the prosody-enhanced stream (from Dec2). This difference is then quantized using a second codebook (VQ2) to create residual tokens.
@@ 2.3 The Structure of the TTS Model
Figure 2: The network structure of the TTS model. The index denotes the order of each token. Best viewed in colors
@@
Autoregressive Generation: The generation process proceeds autoregressively at 12.5Hz. The input to the next step of the primary Semantic Decoder is formed by concatenating its own output feature from the current step with the embeddings of all tokens generated in that step (semantic, prosody, and residual). Generation terminates when a special stop token is predicted by the Semantic Decoder. The speaker embedding is not used during token prediction; it is only provided as a condition to the codec's decoder during the final synthesis stage.
@@ 3.1 Datasets and Preprocessing
The training data of the codec included the English portion of the Multilingual LibriSpeech (MLS) dataset [pratap2020mls], Wenetspeech [zhang2022wenetspeech], Textrolspeech [ji2024textrolspeech], Aishell 1 & 2 [bu2017aishell, du2018aishell], CN-Celeb 1 & 2 [li2022cn], and VoxCeleb 1 & 2  [nagrani2020voxceleb]. The TTS model was trained on MLS-en, Libritts [zen2019libritts], and VCTK [yamagishi2017vctk] only.
All speech audio was resampled to 16kHz. During training, 6-second segments were randomly cropped from each sample. These segments were then transformed into 80-dimensional Mel-spectrograms using a 25ms window length and a 10ms hop length. The codec's reconstruction ability was evaluated on the Librispeech test set, while the zero-shot TTS capability was evaluated on the English test set from Seed-TTS-eval [^1] https://github.com/BytedanceSpeech/seed-tts-eval.
@@ 3.2.1 Codec Configuration
The encoders (Enc1 &\& Enc2) and decoders (Dec2 &\& Dec3) in the codec are convolution-based networks. Their structures are similar to SEANet [tagliasacchi2020seanet], with convolution layer parameters adjusted to match the required frame rates for each stream. Two self-attention[vaswani2017attention] modules are added at the end of the Enc1 and the input of Dec3, respectively, which involve the global relation between frames. Dec1 is a 2-layer conformer network with cross-attention modules, a structure similar to the decoder in CTX-vec2wav [du2024unicats]. A pretrained 16kHz Fre-GAN  [kim2021fre] was used as the vocoder to convert reconstructed Mel-spectrograms into audio waveforms. Three versions of the codec were trained with different bitrates- 424/524/612. For all versions, the semantic stream uses a fixed codebook of 500 centers from the pre-trained Hubert model [^2] https://github.com/facebookresearch/textlesslib/blob/main/textless. The sizes of the codebook for the prosody stream (VQ1) in these three codecs are 64/256/512; their codebook sizes for the residual stream (VQ2) are 32/256/2048, respectively.
@@ 3.2.3 Evaluation Metrics
  • Reconstruction Quality: Measured with Short-Time Objective Intelligibility (STOI) [andersen2017non], Perceptual Evaluation of Speech Quality (PESQ) [rix2001perceptual], and a pre-trained UTMOS [saeki2022utmos] model for naturalness.
  • Speaker Similarity (SIM): Measured between the original and synthesized speech using a WavLM-large-based speaker verification model [chen2022large].
@@ 3.3 Results
Table 1: Comparison of speech reconstruction quality on the Librispeech test set. Our proposed models are benchmarked against a comprehensive set of low-bitrate codecs. `Nq' denotes the number of quantizers and `TPS' is tokens per second. The best scores are in bold.
@@
MSR-Codec-612 500/512/2048 3 62.5 612 0.90 2.73 2.09 4.13 0.83
Table 2: Zero-shot TTS performance on the Seed-TTS English test set. `Data size' is measured in hours. `RTF' denotes Real-Time Factor. Our model's results are bolded to highlight competitive performance from a much smaller model. The best overall score in each column is also marked.
@@
Table 3: Voice conversion results demonstrating information disentanglement. `+S' indicates timbre conversion and `+P' indicates prosody conversion. Key metrics supporting successful and disentangled transfer are in bold.
@@ 3.3.1 Speech Reconstruction Quality
As shown in  Table 1, the proposed codecs achieve strong reconstruction quality at low bitrates. Compared to other codecs, our models demonstrate the highest speaker similarity and comparable UTMOS scores, indicating that timbre and naturalness are well-preserved. The STOI and PESQ scores are relatively lower, which is expected as these metrics are sensitive to signal-level details that require more bits to reconstruct perfectly. Notably, these scores improve significantly when using the larger bitrate version of the codec, where the MSR-Codec-612 achieves the highest STOI and speaker similarity scores.
@@ 3.3.2 Zero-Shot Text-to-Speech
The TTS results in Table 2 highlight our model's exceptional efficiency and performance. As the most lightweight model trained on the least amount of data, our system still achieves a state-of-the-art Word Error Rate (WER), surpassing all competing models. It also delivers the highest speaker similarity, indicating superior voice cloning. Crucially, our model demonstrates its speed by achieving the lowest Real-Time Factor (RTF), making it the fastest generator among the tested systems. This combination of top-tier accuracy and unparalleled efficiency underscores the significant advantages of our proposed architecture.
@@ 3.3.3 Voice and Prosody Conversion
The codec's disentangled representation enables inherent voice conversion (VC) capabilities, which we evaluated to validate the separation of speech attributes. $8$ utterances are randomly sampled from the dataset of VCTK. These utterances come from different speakers (4 male/ 4 female), and they are utilized as the target prompts. $100$ utterances are randomly sampled from the clean test set of Libritts and utilized as the source speeches.
Timbre Conversion: By swapping only the speaker embedding of a source speech with a target one, our method achieves higher target speaker similarity than both CosyVoice2 and Seed-VC [liu2024zero]. As shown in  Table 3, this conversion preserves the original prosody, with the average $F_{0}$ remaining close to the source speech (low $\Delta F_{0,\text{src}}$) and not shifting towards the target (high $\Delta F_{0,\text{tar}}$ ). This confirms the successful disentanglement of the timbre from the prosody.
Prosody Conversion: By keeping the source speaker embedding and semantic tokens but predicting new prosody and residual tokens from a target speech, our model achieves a clear prosody transfer. The results show a significant $F_{0}$ shift towards the target (low $\Delta F_{0,\text{tar}}$) while maintaining the original speaker's timbre.
Combined Conversion: When both the speaker embedding is swapped, and the prosody tokens are re-predicted, the model achieves the best results in transferring both timbre and prosody simultaneously.